### v2

# FDF: Flexible Decoupled Framework for Time Series Forecasting with Conditional Denoising and Polynomial Modeling

Abstract.
Time series forecasting is vital in numerous web applications, influencing critical decision-making across industries. While diffusion models have recently gained increasing popularity for this task, we argue they suffer from a significant drawback: indiscriminate noise addition to the original time series followed by denoising, which can obscure underlying dynamic evolving trend and complicate forecasting. To address this limitation, we propose a novel flexible decoupled framework (FDF) that learns high-quality time series representations for enhanced forecasting performance. A key characteristic of our approach leverages the inherent inductive bias of time series data of its decomposed trend and seasonal components, each modeled separately to enable decoupled analysis and modeling. Specifically, we propose an innovative Conditional Denoising Seasonal Module (CDSM) within the diffusion model, which leverages statistical information from the historical window to conditionally model the complex seasonal component. Notably, we incorporate a Polynomial Trend Module (PTM) to effectively capture the smooth trend component, thereby enhancing the model’s ability to represent temporal dependencies. Extensive experiments validate the effectiveness of our framework, demonstrating superior performance over existing methods and highlighting its flexibility in time series forecasting. The source code is available at https://github.com/zjt-gpu/FDF.
Time Series Forecasting; Diffusion Model; Decoupled Modeling
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Computing methodologies Machine learningccs: Mathematics of computing Time series analysis
Figure 1. The forward diffusion process progressively adds noise to the original time series, causing the trend components to be overwhelmed and obscured, while noise increasingly dominates the series.
1. INTRODUCTION
Time series forecasting is a fundamental task in various domains, playing a pivotal role in driving informed decision-making across industries (Kashpruk et al., 2023; Liu et al., 2024). It involves analyzing data captured in a historical window to model the dynamic evolution of underlying processes and predict future trends over a predicted window (Montgomery et al., 2015; Cheng et al., 2024). In web applications, accurate time series forecasting is essential for tasks such as demand prediction in e-commerce platforms (Trastour et al., 2002), load balancing in cloud computing services (Kilcioglu et al., 2017), and user engagement forecasting (Drutsa et al., 2015) in social media analytics.
Over the years, a multitude of approaches have been proposed to tackle this problem, ranging from classical statistical models like Autoregressive Integrated Moving Average (ARIMA) (Asteriou and Hall, 2011) to advanced deep learning techniques employing recurrent neural networks (RNNs)(Lin et al., 2023; Lai et al., 2018), convolutional neural networks (CNNs)(Bai et al., 2018; Liu et al., 2022; Cheng et al., 2024), and transformer-based networks (Zhou et al., 2021; Wu et al., 2021; Cheng et al., 2023a). Despite their effectiveness, these previous methods often lack consideration for the inherent complexities of time series data, such as non-linear temporal dependencies, continuous value ranges, and the probabilistic nature of future events (Kolambe and Arora, 2024). They may struggle to capture the uncertainty associated with future forecasts or to model complex patterns like heteroscedasticity and non-stationarity. Probabilistic models have emerged as promising alternatives, offering a principled way to quantify uncertainty and model intricate data distributions.
Diffusion models are probabilistic generative methods originally developed for data generation tasks. These models gradually add noise to the data, transforming it into a Gaussian-like distribution, then reverse the process to recover the original data (Ho et al., 2020). Their capacity to model complex stochastic processes and intricate patterns has established them as a powerful tool for time series forecasting. Early models, such as TimeGrad (Rasul et al., 2021), employ autoregressive denoising to predict future states, introducing a novel framework for time series prediction. Recent approaches, such as TimeDiff (Shen and Kwok, 2023), integrate external information to guide the diffusion process, leading to significant improvements in both accuracy and uncertainty quantification. Following these advancements, diffusion models have been successfully applied to tasks such as stock market forecasting (Daiya et al., 2024) and synthetic network traffic generation (Sivaroopan et al., 2024), demonstrating their potential to advance time series forecasting methodologies.
Although their effectiveness, we argue that existing diffusion-based methods in time series forecasting often suffer from a significant drawback: they naively apply noise addition and denoising processes without considering the unique inductive biases inherent in time series data. Specifically, these methods indiscriminately add noise to the original time series and then attempt to denoise it, treating all temporal components equally and ignoring the structured patterns present in the data. As depicted in fig.1, this approach neglects the crucial distinctions between smooth trends and complex seasonal patterns, which are essential for accurate forecasting. As a result, the essential temporal structures may be lost or distorted during the noise addition and removal phases, obscuring underlying dynamic trends and complicating the forecasting task. By disregarding the inductive biases—trend smoothness and seasonal regularities—these methods struggle to effectively model the distinct temporal behaviors inherent in time series data. Addressing this limitation requires a tailored approach that not only preserves critical information during the diffusion process but also explicitly models the unique characteristics of time series components.
To overcome these challenges, we propose a novel flexible decoupled framework(FDF) for time series forecasting that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics. Specifically, we employ meticulously designed conditional denoising seasonal module (CSDM) to capture the complex seasonal component, introducing an innovative denoising strategy conditioned on statistical information derived from the historical window. Concurrently, we model the smooth trend component using polynomial trend module (PTM), which are adept at capturing gradual changes over time. This decoupling method enhances the model’s ability to capture temporal dependencies and complex patterns without sacrificing critical trend information. The key contributions of this work are:
  • We introduce a novel framework that enables the decoupled modeling of trend and seasonal components, providing new insights into time series forecasting.
  • We design tailored methods for each component-using CDSM for the seasonal component and PTM for the trend component—to effectively capture their distinct behaviors.
  • We conduct comprehensive experiments on benchmark datasets to validate the effectiveness of our framework, demonstrating superior performance over existing methods and highlighting its flexibility and effectiveness.
2. RELATED WORK
2.1. Time Series Forecasting
Time series forecasting has evolved from traditional statistical methods to advanced deep learning approaches for more accurate predictions (Cheng et al., 2023b). Early models like VAR (Watson, 1994) and ARIMA (Asteriou and Hall, 2011) were widely used for linear temporal dependencies, but their limitations in capturing complex nonlinear patterns prompted a shift toward deep learning. Recurrent neural networks (RNNs), such as LSTNet (Lai et al., 2018) and SegRNN (Lin et al., 2023), have become popular for modeling sequential dependencies. Concurrently, convolutional neural networks (CNNs) like TCN (Bai et al., 2018) and SCINet (Liu et al., 2022) excel at capturing both local and global temporal features through convolutional filters. In addition, graph-based models such as MTGNN (Wu et al., 2020) and FourierGNN (Yi et al., 2024) have improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models, including Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021), have advanced the field by utilizing self-attention mechanisms to efficiently capture long-range dependencies, proving particularly effective in high-dimensional and long-horizon tasks. Interestingly, simpler linear models like DLinear (Zeng et al., 2023) and TSMixer (Ekambaram et al., 2023) have shown that, by focusing on long-term trends, well-tuned linear layers can outperform more complex architectures, emphasizing efficiency and scalability. While current deep learning methods have strong memory capabilities, they often struggle with modeling complex and volatile fluctuations, presenting challenges in capturing nuanced time series dynamics.
2.2. Diffusion Models
In recent years, Diffusion Denoising Probabilistic Models have advanced time series forecasting by capturing complex temporal patterns. Early models like TimeGrad (Rasul et al., 2021) introduced autoregressive denoising with Langevin sampling to enhance multivariate prediction, while TSDiff (Kollovieh et al., 2024) improved short-term accuracy and data generation through a self-guiding mechanism. Score-based models, such as ScoreGrad (Yan et al., 2021), applied stochastic differential equations (SDEs) for continuous-time forecasting, expanding diffusion into a continuous spectrum and addressing irregularly sampled data. More recently, conditional diffusion models have gained prominence by incorporating external information to guide the diffusion process. For example, TimeDiff (Shen and Kwok, 2023) introduced future mixing and autoregressive initialization, outperforming traditional models in time series prediction. DiffLoad (Wang et al., 2023) enhanced energy load forecasting with uncertainty quantification, while TDSTF (Chang et al., 2023) and DiffECG (Neifar et al., 2023) applied conditional diffusion to improve predictions in sparse ICU data and ECG signals, respectively. Latent Diffusion Models (LDMs) (Rombach et al., 2022) have further improved computational efficiency by performing diffusion in lower-dimensional spaces, as demonstrated by LDCast and Latent Diffusion Transformers (LDT) (Feng et al., 2024), which enhanced precipitation and scalable time series forecasting. DSPD and CSPD (Bilos et al., 2023) extended these approaches to function space for tasks like anomaly detection and interpolation. Additionally, diffusion models have been successfully applied to domains such as flood forecasting (Shao et al., 2024), stock market prediction (Daiya et al., 2024), and EV load forecasting (Li et al., 2024). Despite the promising results of diffusion models in time series forecasting, they lack effective methods for separating and modeling trend components, limiting their ability to capture long-term dependencies and trend shifts.
Figure 2. Illustration of the proposed Flexible Decoupled Framework (FDF). The Polynomial Trend Module (PTM) models the trend component to capture different patterns, while the Conditional Denoising Seasonal Module (CDSM) refines the seasonal component through a diffusion process conditioned on historical data.
3. METHODOLOGY
In this section, we formalize the time series forecasting problem and present the proposed Flexible Decoupled Framework (FDF), comprising two core components: the Conditional Denoising Seasonal Module (CDSM) and the Polynomial Trend Module (PTM).
3.1. Problem Statement
Time series represent sequences of data points ordered chronologically, typically recorded at regular intervals. Formally, consider a multivariate time series $X_{1:L+T}\in\mathbb{R}^{(L+T)\times d}$, where $L$ denotes the length of the historical window and $T$ denotes the length of the predicted window, with $d$ attributes recorded for each time step. The objective of time series forecasting is to train a model $G$ that learns to map the historical series $X_{L}$ to the predicted series $X_{T}$. The model leverages the information embedded in the historical window to predict future values, aiming to capture both temporal dependencies for accurate forecasting.
3.2. Overall Framework
The Flexible Decoupled Framework (FDF) addresses time series forecasting by independently modeling trend and seasonal patterns, thereby isolating and capturing distinct temporal behaviors more effectively. As depicted in Fig. 2, FDF consists of two main modules: the CDSM and the PTM. The CDSM employs a denoising diffusion process conditioned on statistical information from the historical window to accurately capture and refine seasonal variation patterns. In parallel, the PTM models trends using polynomial linear layers, providing flexibility in handling both linear and nonlinear trends. By decoupling and specifically modeling these components, FDF enhances forecasting accuracy, offering an effective solution for capturing the complex temporal dynamics inherent in time series. By handling trend and seasonal components independently, FDF allows each module to focus on its respective dynamics, leading to enhanced overall performance in forecasting.
3.3. Decoupling Preprocessing
3.3.1. Instance Normalization
Consistent normalization across historical and predicted windows is critical for FDF framework stability. We standardize the time series in the historical window by subtracting the mean and dividing by the standard deviation (Kim et al., 2021). The same mean and standard deviation are transferred to the predicted window during training. Specifically, $x_{L}$ is normalized to $x_{L}^{0}$, and $x_{T}$ is normalized to $x_{T}^{0}$ using the statistics from $x_{L}$. This ensures the predicted window is aligned with the historical window, preserving temporal structure and stability throughout the forecasting process.
3.3.2. Decomposition Approach
Decoupling is a fundamental aspect of time series analysis, making it natural to decompose the series into its underlying trend and seasonal components. This enables the model to handle each temporal characteristic independently, ensuring that the distinct behaviors of different components are effectively captured (Cleveland et al., 1990). We decompose both historical $X_{L}^{0}$ and predicted $X_{T}^{0}$ windows into trend and seasonal components. The trend component for the historical window, $X_{L,t}^{0}$, is estimated using a moving average:
(1) $X_{L,t}^{0}=\text{Avg}(X_{L}^{0}),$
where Avg is the moving average of the historical series $X_{L}^{0}$.
The seasonal component, $X_{L,s}^{0}$, is then obtained by subtracting the trend from the original series:
(2) $X_{L,s}^{0}=X_{L}^{0}-X_{L,t}^{0}.$
Similarly, for the predicted window, we calculate $X_{T,t}^{0}$ and $X_{T,s}^{0}$:
(3) $X_{T,t}^{0}=\text{Avg}(X_{T}^{0}),$
(4) $X_{T,s}^{0}=X_{T}^{0}-X_{T,t}^{0}.$
This decomposition method enhancing accuracy by focusing on each component’s characteristics.
3.4. Conditional Denoising Seasonal Module
Modeling complex seasonal components in time series poses significant challenges due to their intricate patterns and non-stationary behaviors. Therefore, we propose the CDSM, which employs a denoising diffusion model specifically designed to model the seasonal component. By incorporating historical statistical information as conditional inputs, the CDSM better captures seasonal patterns. This conditional integration enables the model to handle intricate temporal dynamics more effectively, aligning seasonal predictions with historical contexts and improving forecasting accuracy.
3.4.1. Forward Process
In the forward diffusion process, we simulate the degradation of the seasonal signal by incrementally adding noise to the seasonal component of the target window. At each diffusion step $k$, the noisy seasonal series is formulated as:
(5) $X_{T,s}^{k}=\sqrt{\bar{\alpha}_{k}}\,x_{T,s}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon,$
where $X_{T,s}^{k}$ represents the seasonal component at step $k$, $x_{T,s}^{0}$ is the original seasonal component, $\epsilon\sim\mathcal{N}(0,I)$ is standard Gaussian noise, and $\bar{\alpha}_{k}\in[0,1]$ is determined by the noise schedule. As $k$ increases, the seasonal component progressively transitions towards a noise-dominated state.
3.4.2. Conditioned Denoising Network
The denoising process starts by embedding $X_{T,s}^{k}$ into a hidden space:
(6) $\mathbf{h}=\text{Emb}(X_{T,s}^{k}),$
where $\text{Emb}(\cdot)$ is a linear projection that transforms the noisy sequence into a hidden dimension $d_{\text{model}}$. To provide the model with diffusion step-specific information, sinusoidal positional encodings (Ho et al., 2020; Gu et al., 2022) are applied to encode the diffusion step $t$. These encodings are integrated into the network via Adaptive Layer Normalization (AdaLN):
(7) $\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t},$
where $\alpha_{t}$ and $\beta_{t}$ are learnable parameters projected from the diffusion step encoding. This mechanism allows the network to adjust its behavior dynamically, addressing the challenge of effectively utilizing time step information in the denoising process.
The decoder captures both global temporal dependencies and feature interactions by processing the embedded features through a combination of convolutional layers and multilayer perceptrons (MLPs). The convolutional layers specialize in extracting fine-grained local temporal variations, while the MLP layers focus on modeling cross-dimensional interactions by capturing relationships across various dimensions. Finally, layer normalization and a projection layer are applied to map the model dimension $d_{\text{model}}$, back to the original feature space of dimension $d$.
To enhance the denoising capability, we incorporate conditional information based on the historical seasonal component $X_{L,s}^{0}$ (Pierce, 1978). The historical data is segmented into patches $P$ of length $H$, with each patch capturing localized segments of the seasonal signal. For each patch $P_{L,i}$, we compute the local mean $\mu_{L,i}$ and variance $\sigma_{L,i}^{2}$ to represent its statistical properties:
(8) $\mu_{L,i}=\frac{1}{H}\sum_{h=1}^{H}x_{L,s,i,h},$
(9) $\sigma_{L,i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{L,i}\right)^{2}.$
These local statistics capture variability within the seasonal component, adapting to localized changes in the seasonal patterns. We then predict the corresponding mean $\hat{\mu}_{T,i}$ and variance $\hat{\sigma}_{T,i}^{2}$ for the predicted window $P_{T,i}$ using two independent MLPs:
(10) $\hat{\mu}_{T,i}=\text{MLP}_{\mu}(\mu_{L,i}),$
(11) $\hat{\sigma}_{T,i}^{2}=\text{MLP}_{\sigma}(\sigma_{L,i}^{2}).$
The predicted statistics are integrated into the denoising process through Gaussian sampling:
(12) $\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{T,i}+\hat{\sigma}_{T,i}\cdot z,$
where $z\sim\mathcal{N}(0,I)$. The term $\nabla_{\mathbf{x}_{T,s}}^{k-1}$ represents the conditional information, and sampling $z$ from a normal distribution using the predicted mean and variance allows the model to incorporate stochastic variability in the seasonal component. By leveraging the statistical properties of the historical window to predict future values, the model ensures that its predictions align with historical seasonal patterns, thereby enhancing predictive accuracy.
Finally, the denoised output $\hat{x}_{T,s}^{0}$ is combined with the conditional information through a weighted summation:
(13) $X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1},$
where $\rho_{1}$ and $\rho_{2}$ are learnable parameters that balance the contributions of the denoised series and the conditional information. This combination ensures that historical seasonal patterns are effectively captured while preserving statistical characteristics, enabling the model to handle complex temporal variations.
3.5. Polynomial Trend Module
Different trends exhibit distinct patterns of progression, including periods of growth and decline, making it essential to capture a broader range of trend behaviors. By leveraging Cover’s theorem (Cover, 1965), we project complex time series patterns into a higher-dimensional space, enabling the extraction of both linear and nonlinear trends. Based on this, polynomial modeling (Liu et al., 2021) is employed to better adapt to these diverse trend variations. To fully harness the potential of polynomial modeling, we introduce a square root transformation as a specific form of nonlinear transformation within the polynomial framework. The square root transformation effectively smooths extreme values, enhancing the model’s flexibility in capturing a wide range of trend patterns.
The PTM operates through two complementary pathways. The first pathway models stable trends by applying a linear layer directly to the historical trend series:
(14) $T_{\text{origin}}(x_{T,t}^{0})=\text{Linear}(x_{T,t}^{0}),$
where the Linear layer captures the overall shape of the trend and adapts to various temporal structures.
The second pathway enhances trend modeling by applying a square root transformation to the trend component, thereby mitigating the impact of extreme values and stabilizing trend predictions, particularly in the presence of noise:
(15) $T_{\text{sqrt}}(x_{T,t}^{0})=\text{Linear}(\sqrt{x_{T,t}^{0}}).$
By applying the linear layer after the root transformation, the model better captures diverse trend patterns.
The outputs from both pathways are combined through a weighted summation to capture a broader range of trends:
(16) $f_{\theta}(x_{T,t}^{0})=\lambda_{1}T_{\text{origin}}(x_{T,t}^{0})+\lambda_{2}T_{\text{sqrt}}(x_{T,t}^{0}),$
where $\lambda_{1}$ and $\lambda_{2}$ are learnable parameters that dynamically modulate the influence of the linear and root-based components, optimizing the representation of trend dynamics. This flexible mechanism allows the model to seamlessly adapt between linear and sub-linear growth patterns, ultimately enhancing its ability to accurately forecast complex time series that exhibit diverse temporal dynamics.
Algorithm 1 FDF Training Algorithm.
1:  Require: Number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$, initialized model parameters $\theta$.
2:  repeat
3:     Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, obtaining $x_{L}^{0}$; apply the same normalization to $x_{T}$, yielding $x_{T}^{0}$.
4:     Decompose $x_{L}^{0}$ into trend $x_{L,t}^{0}$ and seasonal $x_{L,s}^{0}$ components using (2) and decompose $x_{T}^{0}$ similarly using (4).
5:     Compute the cumulative product of the noise schedule: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
6:     Sample a diffusion step $k$, and generate $x_{T,s}^{k}$ using (5).
7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ via (12).
8:     Use the predicted conditional information to estimate the seasonal component via (13).
9:     Predict the trend component according to (16).
10:     Integrate the predicted components using (17): $\hat{x}_{T}^{0}$.
11:     Denormalize $\hat{x}_{T}^{0}$ using $\mu_{L}$ and $\sigma_{L}^{2}$, obtaining $\hat{x}_{T}$.
12:     Compute the loss $L_{\text{mse}}(\theta)$ as defined in (18).
13:     Update the model parameters by taking a gradient descent step on $\nabla_{\theta}L_{\text{mse}}$.
14:  until convergence.
Algorithm 2 FDF Inference Algorithm.
1:  Require: Model parameters $\theta$, number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$.
2:  Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, yielding $x_{L}^{0}$, then decompose by (2) and (4).
3:  Initialize $x_{T,s}^{K}\sim\mathcal{N}(0,I)$.
4:  for $k=K$ to $1$ do
5:     Sample noise $\epsilon\sim\mathcal{N}(0,I)$ if $k>1$; otherwise, set $\epsilon=0$.
6:     Compute the cumulative product: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ using (12).
8:     Predict the seasonal component using formula (13).
9:     Update the denoised estimate:
$\hat{x}_{T,s}^{k-1}=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}x_{T,s}^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{x}_{T,s}^{0}+\sigma_{k}\epsilon.$
10:  end for
11:  Predict the trend component using formula (16).
12:  Integrate the trend and seasonal components, denormalize, and obtain the final output $\hat{x}_{T}$.
13:  Return $\hat{x}_{T}$.
3.6. Fusion and Optimization
The integration of trend and seasonal components is essential for enhancing forecasting accuracy. By combining these two aspects, the framework leverages the strengths of both trend and seasonal modeling, resulting in more accurate and comprehensive predictions. Additionally, the denormalization process ensures that the forecasted values are restored to their original distribution, maintaining consistency with the input data. Finally, applying the derived loss function enables joint optimization during training, allowing both components to be learned concurrently and effectively.
3.6.1. Fusion
Algorithms 1 and 2 illustrate the training and inference processes. During both phases, after the trend component $\hat{x}_{T,t}^{0}$ is predicted by the PTM and the seasonal component $\hat{x}_{T,s}^{0}$ is recovered by the CDSM.
The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
(17) $\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$
where $X_{\theta}$ represents the CDSM and $f_{\theta}$ corresponds to the PTM.
3.6.2. Optimization
The final prediction is subsequently denormalized $\hat{x}_{T}^{0}$ using the mean and variance derived from the historical window to $\hat{x}_{T}$, ensuring that the forecast remains consistent with the original distribution of the time series.
The loss function in FDF is derived from the evidence lower bound (ELBO), commonly employed in diffusion models, to minimize the discrepancy between the actual time series $x_{T}^{0}$ and the predicted series $\hat{x}_{T}^{0}$. This formulation ensures that both the trend and seasonal components effectively contribute to the final forecast. The loss function is given by:
(18) $L_{\text{elbo}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2}.$
Additional information on the formulation of the loss function can be found in the appendix B.
4. EXPERIMENTS
Table 1. Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
Dataset ETTh1 ETTm1 Wind Exchange Weather Electricity
Dimension 7 7 7 8 21 321
Frequency 1 hour 15 mins 15 mins 1 day 10 mins 1 hour
T 168 192 192 14 672 168
Table 2. The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
Dataset T FDF Patchmixer TSMixer PatchTST Dlinear Autoformer Diffusion-TS D3VAE CSDI
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 168 0.4364 0.4363 0.4610 0.4560 0.4418 0.4401 0.4401 0.4463 0.4372 0.4420 0.6046 0.5634 1.6559 1.0400 0.9523 0.8618 1.1090 0.8009
ETTh2 168 0.3393 0.3826 0.3731 0.4055 0.3654 0.3981 0.3805 0.4036 0.3976 0.4233 0.4437 0.4658 3.2225 1.4736 5.3165 1.9796 2.0700 1.0748
ETTm1 192 0.3532 0.3781 0.3731 0.3983 0.3673 0.3924 0.3626 0.3890 0.3706 0.3968 0.6299 0.5373 1.6769 1.0326 0.6341 0.6833 1.0770 0.7788
ETTm2 192 0.2303 0.3009 0.2643 0.3226 0.2495 0.3108 0.2472 0.3129 0.2498 0.3224 0.3259 0.3755 2.9380 1.3732 4.1001 1.5884 1.6051 0.9397
Electricity 168 0.1606 0.2526 0.1559 0.2505 0.1563 0.2497 0.1537 0.2494 0.1628 0.2581 0.4125 0.4671 1.4888 0.9935 1.7061 0.9273 —— ——
Exchange 14 0.0204 0.0979 0.0430 0.1150 0.0209 0.0955 0.0293 0.1178 0.0412 0.1320 0.5960 0.5251 2.7192 1.3573 4.8311 1.9942 1.3023 0.8719
Weather 672 0.3292 0.3373 0.3295 0.3411 0.3316 0.3430 0.3299 0.3400 0.3398 0.3801 0.3851 0.3976 —— —— —— —— 0.4875 0.4473
Wind 192 1.1262 0.7534 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 1.3222 0.8489 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
4.1. Experimental Settings
4.1.1. Datasets
We evaluate the effectiveness of our proposed framework on multiple real-world time series datasets from various domains. The datasets include ETTh and ETTm (Zhou et al., 2021), which represent electricity transformer temperature data; Exchange, which contains exchange rates from eight countries; and Weather [^1] https://www.bgc-jena.mpg.de/wetter/, consisting of 21 meteorological indicators sampled at 10-minute intervals. Additionally, we use the Electricity dataset [^2] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which records electricity consumption from 321 clients, and Wind (Li et al., 2022), comprising wind power measurements sampled every 15 minutes from 2020 to 2021.
To ensure consistency in chronological data splitting, we adopt a 6:2:2 split for the ETTh1 and ETTm1 datasets, while employing a 7:1:2 split for the Wind, Weather, Electricity, and Exchange datasets. This setup allows for thorough testing of the model’s generalization across different temporal segments. In alignment with TimeDiff (Shen and Kwok, 2023), we evaluate performance using historical window lengths of ${96,192,720,1440}$ to accommodate varying forecasting horizons. Further details regarding datasets and experimental protocols are provided in Appendix A.2.
4.1.2. Baselines
We compare our proposed FDF model against a diverse set of baseline models representing various methodological approaches.
  • Diffusion models: CSDI (Tashiro et al., 2021), Diffusion-TS (Yuan and Qiao, 2024).
  • Transformer-based models: PatchTST (Nie et al., 2023), Autoformer (Wu et al., 2021).
  • Linear-based models: DLinear (Zeng et al., 2023), TSMixer (Ekambaram et al., 2023).
  • Hybrid model: PatchMixer (Gong et al., 2023), D3VAE (Li et al., 2022).
More details of baseline models are available in the Appendix A.3.
4.1.3. Implementation Details
Our proposed FDF model is implemented using PyTorch and trained with an Exponential Learning Rate Scheduler (Paszke et al., 2019), starting with an initial learning rate of $1\times 10^{-3}$ and a batch size of 16. Early stopping with a patience of 10 epochs is employed to prevent overfitting. The diffusion process utilizes $K=50$ steps with a cosine variance schedule (Rasul et al., 2021) from $\beta_{1}=10^{-4}$ to $\beta_{K}=0.5$. We adopt the DDIM sampler (Song et al., 2021). All experiments are conducted on a workstation equipped with an NVIDIA GeForce RTX 4090 GPU. Additional implementation details and hyperparameter settings are provided in Appendix A.4.
4.2. Experiment Results
4.2.1. Main results
We assess the performance of the Flexible Decoupled Framework (FDF) across several real-world datasets, as summarized in Table 2. FDF consistently outperforms state-of-the-art models, demonstrating notable improvements in Mean Squared Error (MSE) and Mean Absolute Error (MAE) over methods such as Diffusion-TS, D3VAE, and PatchMixer, confirming its effectiveness for time series forecasting. On challenging datasets like ETTh and ETTm, FDF adeptly manages varying time resolutions and intensities. In the Weather dataset, FDF sustains high accuracy across extended forecasting windows, underscoring its strength in forecasting tasks. While the Electricity dataset presents challenges due to its higher dimensionality, FDF still maintains competitive performance. The Conditional Denoising Seasonal Module (CDSM) captures complex fluctuations, while the Polynomial Trend Module (PTM) effectively preserves trends, capturing both linear and nonlinear dynamics. Together, these modules enhance FDF’s adaptability across diverse scenarios, a capability that is also validated by the more detailed experimental results in Appendix C.1.
Figure 3. Visual comparison of CDSM and PTM in modeling trend components.
Figure 4. Visual comparison of CDSM and PTM in modeling seasonal components.
Figure 5. Parameter sensitivity analysis of FDF.
4.2.2. Component Module Experiment
In this section, we analyze the FDF model’s performance in modeling trend and seasonal components by comparing CDSM and PTM on the ETTh1 dataset with an input length of 192 and a prediction length of 168. The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy. As shown in Figure 3, PTM excels in capturing smooth and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, characterized by complex fluctuations, CDSM outperforms PTM, as shown in Figure 4. CDSM effectively captures sharp seasonal changes, while PTM smooths out critical variations, focusing on underlying regularities in the fluctuations, but sometimes missing rapid shifts. These results validate each module’s modeling capabilities for its designated component. By assigning each module to the component it best models, FDF captures both distinct trends and complex fluctuations, enhancing forecasting accuracy, as seen in both trend and seasonal visualizations.
Figure 6. Evaluation of decoupling variants.
4.2.3. Decoupling Variants Experiments
We evaluated the impact of the decoupling strategy on the performance of our proposed FDF model using the ETTh2 and ETTm2 datasets. As shown in Figure 6, we compared the following two model variants:
  • FDFw/o ${}_{\text{w/o decoupling}}$: This variant bypasses the decoupling process, directly applying the CDSM to the original time series.
[…]
4.2.4. Ablation Experiments
Table 3. Ablation study results on ETTh2 and ETTm2.
Method ETTh2 ETTm2
MSE MAE MSE MAE
FDF 0.3393 0.3826 0.2303 0.3009
FDFw/o ${}_{\text{w/o CR}}$ 0.3404 0.3781 0.2320 0.3023
[…]
FDFw/o ${}_{\text{w/o INR}}$ 0.3853 0.4215 0.2459 0.3272
FDFw/o ${}_{\text{w/o CR+PLR}}$ 0.3442 0.3807 0.2335 0.3015
FDFw/o ${}_{\text{w/o INA}}$ 0.2300 0.3006 0.2300 0.3006
We present the results of ablation experiments assessing the contributions of individual components in the FDF architecture (Table 3):
  • FDFw/o ${}_{\text{w/o CR}}$: Removes the conditional information.
[…]
  • FDFw/o ${}_{\text{w/o INR}}$: Removes instance normalization during both the training and inference phases.
  • FDFw/o ${}_{\text{w/o CR+PLR}}$: Removes the conditional information and replaces PTM with a linear layer.
  • FDFw/o ${}_{\text{w/o INA}}$: Removes instance normalization alignment.
The ablation results emphasize the critical role of each component in the FDF architecture, particularly in handling complex temporal dynamics across datasets such as ETTh2 and ETTm2. Removing the condition information injection module leads to a performance decrease, suggesting that while the module aids in incorporating external factors, its absence has a smaller impact, especially in coarser datasets like ETTh2. In contrast, replacing the PTM with a simple linear layer causes a more substantial performance drop, underscoring the importance of non-linear trend modeling, particularly in finer-grained datasets like ETTm2. Eliminating instance normalization results in the largest decline, particularly in ETTm2, highlighting its role in stabilizing training for high-resolution datasets with frequent fluctuations. The combined removal of the condition information injection module and PTM replacement further degrades performance, illustrating the necessity of both components. Interestingly, removing instance normalization during training improves performance on ETTh2, suggesting it is less critical in datasets with clearer trend-seasonal separation. Overall, these findings demonstrate the adaptability of the FDF architecture, with each component’s contribution varying according to the dataset’s granularity. Further details are in Appendix C.3.
4.2.5. Hyper-parameter sensitiviy analysis
We conducted a hyperparameter sensitivity analysis on the ETTh1 dataset to explore the impact of key design choices in FDF—hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon—on its performance. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, but larger values such as $d_{\text{model}}=512$ lead to diminishing returns and potential overfitting, indicating the need to balance model capacity and computational efficiency. Similarly, $H$ significantly affects performance, with $H=8$ effectively capturing both local dependencies and global temporal patterns, while larger values introduce redundancy and reduce efficiency. Finally, FDF maintains consistent performance across different prediction horizons, with the PTM capturing trends and the diffusion-based seasonal component adapting to fluctuations, ensuring reliable forecasting. Overall, the analysis confirms FDF’s balance between complexity and scalability, making it suitable for a wide range of forecasting tasks.
Figure 7. Visualization of trend and seasonal components for the electricity dataset.
4.2.6. PTM & CDSM Joint Modeling Capability Visualization
Figure 7 shows the combined performance of the PTM and CDSM sub-modules in the FDF framework. By comparing predicted and actual trend and seasonal components, the framework’s interpretability and modeling capabilities are validated. The FDF effectively decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its strength in modeling complex temporal dynamics. In Figure 7 (a), applied to the Electricity dataset with clear periodic and trend characteristics, PTM captures smooth, stable trends with predictions closely matching actual trends, thanks to its ability to model the overall temporal structure. Figure 7 (b) showcases CDSM’s ability to capture complex seasonal variations. The predicted seasonal components align well with actual patterns, compensating for PTM’s trend deviations and further improving accuracy. This highlights CDSM’s flexibility in handling noisy data with intricate cyclic features. Overall, the joint modeling of PTM and CDSM within the FDF framework demonstrates strong complementarity, enabling enhanced interpretability and predictive accuracy in datasets with both trend and seasonal components.
5. Conclusion
In this paper, we introduce the Flexible Decoupled Framework (FDF), a novel architecture for time series forecasting that decouples trend and seasonal components using specialized modules tailored for each. This decoupling strategy enables more precise modeling of temporal dynamics, with the Polynomial Trend Module (PTM) capturing both smooth and complex trends, and the Conditional Denoising Seasonal Module (CDSM) effectively addressing intricate seasonal fluctuations. Extensive experiments on multiple real-world datasets demonstrate that FDF consistently outperforms state-of-the-art models, achieving significantly lower MSE and MAE scores. The modular design, which separately handles distinct temporal patterns, represents a substantial advancement in time series forecasting, offering improved efficiency and accuracy for complex forecasting tasks. Future work will explore extending the framework to handle irregular time series and expanding its application to broader domains and scenarios.
Appendices
Appendix A Experiment Details
A.1. Denoising Network Details
The proposed FDF framework consists of several key subnetworks: the feature embedding network, the denoising decoder network, and the projector module. Table 4 details the input and output sizes of each subnetwork, where $d$ is the number of variables in the time series, $L$ is the sequence length, and $H$ is the hidden dimension.
Table 4. Input and output sizes of the denoising networks.
Network Input size Output size
Embedding $T\times d$ $T\times d_{\text{model}}$
Denoising decoder $T\times d_{\text{model}}$ $T\times d_{\text{model}}$
Projector $T\times d_{\text{model}}$ $T\times d$
The key parameters corresponding to each dataset are shown in Table 5. $d_{\text{hidden}}$ represents the hidden dimension used during the conditional information injection process, $d_{\text{model}}$ denotes the hidden dimension within the denoising network, $H$ indicates the patch size, and $LR$ refers to the learning rate.
Table 5. Key parameters corresponding to each dataset.
Dataset $d_{\text{model}}$ $H$ $LR$
ETTh 256 8 0.001
ETTm 256 4 0.001
Exchange_rate 256 4 0.05
Weather 512 4 0.008
Wind 256 4 0.008
Electricity 1024 4 0.002
A.2. Datasets
We perform extensive experiments on several real-world time series datasets across diverse domains to assess the performance of our proposed framework. The datasets used in this study are: (i) ETTh and ETTm (Zhou et al., 2021), comprising two years of electricity transformer temperature data from China, recorded at 1-hour and 15-minute intervals, respectively. These datasets offer valuable insights into various trends and complex variations in electricity usage. (ii) Exchange (Lai et al., 2018), which includes daily exchange rates from eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore. This dataset presents challenges due to the complex dependencies and volatility inherent to financial time series. (iii) Weather [^3] https://www.bgc-jena.mpg.de/wetter/, containing 21 meteorological indicators recorded every 10 minutes from 2020 to 2021, providing data that captures seasonal patterns and sudden weather changes. (iv) Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which provides hourly electricity consumption data from 321 clients over two years. This dataset is characterized by complex behaviors influenced by external factors. (v) Wind (Li et al., 2022), containing wind power generation data recorded at 15-minute intervals from 2020 to 2021, reflecting dynamics shaped by daily routines and unforeseen events.
We split the ETTh and ETTm datasets into training, validation, and test sets using a 6:2:2 ratio, and employ a 7:1:2 split for the Weather, Electricity, and Exchange datasets. These splits ensure chronological consistency, crucial for realistic forecasting and preventing data leakage. Following TimeDiff (Shen and Kwok, 2023), we select historical context lengths from ${96,192,720,1440}$, allowing for the evaluation of model performance across various forecasting horizons.
A.3. Baselines
To comprehensively assess the effectiveness of our proposed FDF framework, we compare it against a diverse set of baseline models, each representing distinct methodological approaches:
A.4. Implementation Details
Our proposed TTDN model is implemented using PyTorch and trained with an Exponential Learning Rate Scheduler (Paszke et al., 2019) to dynamically adjust the learning rate during training. The initial learning rate is set to $1\times 10^{-3}$ and decays exponentially to ensure smoother convergence. We employ a batch size of 16 and apply early stopping with a patience of 10 epochs to prevent overfitting. The diffusion process in TTDN uses $K=50$ diffusion steps, following a cosine variance schedule (Rasul et al., 2021), starting from $\beta_{1}=10^{-4}$ and increasing to $\beta_{K}=0.5$. This schedule balances the trade-off between noise addition and signal preservation during diffusion.
Several efficient learning-free samplers have been introduced to accelerate inference in diffusion models, such as DDIM (Song et al., 2021), Analytic-DPM (Bao et al., 2022), and DPM-Solver (Lu et al., 2022). To expedite inference in our model, we utilize the Denoising Diffusion Implicit Models (DDIM) sampler (Song et al., 2021), which reduces the number of diffusion steps required for sampling without compromising performance. This is particularly advantageous for real-time forecasting applications, where computational efficiency is paramount.
For model evaluation, we use the Mean Squared Error (MSE) to measure prediction accuracy. The MSE is defined as:
(19) $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2},$
where $\hat{y}_{i}$ is the predicted value, $y_{i}$ is the true value, and $n$ is the number of observations. Lower MSE values indicate better predictive performance. Additionally, the baseline loss is defined as:
(20) $L_{\text{mse}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2},$
providing a reliable metric for optimization.
To account for the stochastic nature of diffusion processes, we average prediction results over 10 iterations for each time series model. This approach provides an effective estimate of the model’s performance by mitigating random fluctuations in predictions.
The experiments are conducted on a workstation with an NVIDIA GeForce RTX 4090 GPU (24 GB memory). This hardware setup ensures that computational resources do not limit model performance, enabling efficient training and inference.
Appendix B DERIVED ELBO for FDF
We begin by considering the standard diffusion model, which involves a forward process where data is gradually corrupted by noise, and a backward process to denoise it.
In the forward diffusion process, the input $x^{0}$ is transformed into a noisy vector $x^{k}$ through the following steps:
(21) $q(x^{k}|x^{k-1})$ $=\mathcal{N}\left(x^{k};\sqrt{1-\beta_{k}}\,x^{k-1},\beta_{k}\,I\right),$
where $\beta_{k}$ is the variance schedule, and $I$ is the identity matrix.
By composing these steps, we can express the distribution of $x^{k}$ given $x^{0}$ as:
(22) $q(x^{k}|x^{0})$ $=\mathcal{N}\left(x^{k};\sqrt{\bar{\alpha}_{k}}\,x^{0},\left(1-\bar{\alpha}_{k}\right)I\right),$
where $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
Thus, $x^{k}$ can be directly expressed as:
(23) $x^{k}$ $=\sqrt{\bar{\alpha}_{k}}\,x^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$
The backward denoising process is defined by a neural network that predicts the mean and variance:
(24) $p_{\theta}(x^{k-1}|x^{k})$ $=\mathcal{N}\left(x^{k-1};\mu_{\theta}(x^{k},k),\Sigma_{\theta}(x^{k},k)\right),$
where $\Sigma_{\theta}(x^{k},k)$ is fixed as $\sigma_{k}^{2}I$, and $\mu_{\theta}(x^{k},k)$ is modeled by a neural network.
The training objective minimizes the KL divergence between the forward and reverse processes:
(25) $L_{k}$ $=D_{\text{KL}}\left(q(x^{k-1}|x^{k},x^{0})\,\|\,p_{\theta}(x^{k-1}|x^{k})\right).$
For stable training, $q(x^{k-1}|x^{k},x^{0})$ is simplified as:
(26) $q(x^{k-1}|x^{k},x^{0})$ $=\mathcal{N}\left(x^{k-1};\mu_{k}(x^{0},x^{k}),\tilde{\beta}_{k}I\right),$
where $\mu_{k}(x^{0},x^{k})$ is expressed as:
(27) $\mu_{k}(x^{0},x^{k})$ $=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x^{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}.$
The training objective simplifies to:
(28) $L_{k}$ $=\frac{1}{2\sigma_{k}^{2}}\,\mathbb{E}_{k}\left[\left\|\mu_{k}(x^{0},x^{k})-\mu_{\theta}(x^{k},k)\right\|^{2}\right].$
The mean of $x^{k-1}$, denoted $\mu_{x}(x_{\theta})$, is calculated as:
(29) $\mu_{x}(x_{\theta})$ $=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x_{\theta}(x^{k},k).$
The corresponding simplified loss is:
(30) $L_{x}$ $=\omega\mathbb{E}_{k,x^{0},\epsilon}\left[\left\|x^{0}-x_{\theta}(x^{k},k)\right\|^{2}\right],$
where the parameter $\omega$ is a constant, and during training, it is typically set to 1 (Ho et al., 2020).
Now , we introduce normalize the $x_{L}$ and $x_{T}$, and then decompose the condition window $x_{L}^{0}$ and predicted series $x_{T}^{0}$ into its trend and seasonal components:
(31) $x_{L}^{0}$ $=x_{L,t}^{0}+x_{L,s}^{0}.$
(32) $x_{T}^{0}$ $=x_{T,t}^{0}+x_{T,s}^{0}.$
We predict the trend component $x_{t}^{0}$ using $f_{\theta}$:
(33) $\hat{x}_{T,t}^{0}$ $=f_{\theta}(x_{L,t}^{0}).$
Next, the seasonal component $x_{s}^{0}$ undergoes the diffusion process. The noisy version of the seasonal component, denoted by $r_{s}^{k}$, is obtained by adding noise to $x_{s}^{0}$:
(34) $x_{T,s}^{k}$ $=\sqrt{\bar{\alpha}_{k}}x_{T,s}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon.$
We improve the denoising process by integrating local statistical features derived from the historical sequence $x_{L,s}^{0}$. The sequence is segmented into patches $P$, each with a length $H$, to capture complex seasonal variations. For each patch $P_{i}$, we compute the mean $\mu_{i}$ and variance $\sigma_{i}^{2}$, which encapsulate the local statistical properties:
(35) $\mu_{i}=\frac{1}{H}\sum_{h=1}^{H}x_{L,s,i,h}.$
(36) $\quad\sigma_{i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{i}\right)^{2}.$
These computed statistics are utilized by a multi-layer perceptron (MLP) to predict the mean $\hat{\mu}_{i+1}$ and variance $\hat{\sigma}_{i+1}^{2}$ for the next patch:
(37) $\hat{\mu}_{i+1}=\text{MLP}(\mu_{i}).$
(38) $\quad\hat{\sigma}_{i+1}^{2}=\text{MLP}(\sigma_{i}^{2}).$
The conditional input for denoising is then sampled as:
(39) $\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{i+1}+\hat{\sigma}_{i+1}\cdot z,\quad z\sim\mathcal{N}(0,1).$
This allows the seasonal component $\hat{x}_{T,s}^{0}$ to be predicted as:
(40) $\hat{x}_{T,s}^{0}=X_{\theta}(x_{T,s}^{k},k\mid\nabla_{\mathbf{x}_{T,s}}^{k}).$
The final reconstruction of the time series is obtained by combining the predicted trend and seasonal components:
(41) $\hat{x}_{T}^{0}$ $=\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}.$
To train the model, we minimize the mean squared error (MSE) between the true and predicted values for both components. The individual losses are defined as follows:
The trend loss measures the error between the true and predicted trend components:
(42) $L_{\text{trend}}$ $=\mathbb{E}_{x_{T}^{0}}\left[\left\|x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right\|^{2}\right].$
The seasonal loss captures the error between the true and predicted seasonal components:
(43) $L_{\text{seasonal}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left\|x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right\|^{2}\right].$
We introduce a cross-term loss to account for the interaction between the trend and seasonal components. This term ensures the model considers any dependency between these components during prediction.
The cross-term loss is defined as:
(44) $L_{\text{cross}}$ $=2\mathbb{E}_{x_{T}^{0}}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\mathbb{E}_{\epsilon}\left[\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\mid x_{T}^{0}\right]\right]$
(45) $=2\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\right].$
The total loss function is the combination of the trend, seasonal, and cross-term losses:
(46) $L_{\text{mse}}$ $=L_{\text{trend}}+L_{\text{seasonal}}+L_{\text{cross}}.$
Expanding the total loss, we obtain:
(47) $L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\left(\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}\right)\right\|^{2}.$
Finally, we denormalize and integrate the diffusion model into the seasonal component, the loss function becomes:
(48) $L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2}.$
This total loss function balances the contributions of the trend and seasonal components while accounting for their interactions. By minimizing this loss, the model learns to reconstruct the time series with greater accuracy.
Appendix C EXPERIMENTAL RESULTS
C.1. Main Results
Table 6. The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
Dataset L T Diffts Patchmixer TSMixer PatchTST DLinear Autoformer Diffusion-TS D3VAE CSDI
mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
ETTh1 96 168 0.4329 0.4247 0.4337 0.4309 0.4224 0.4182 0.4300 0.4304 0.4325 0.4326 0.5336 0.5038 1.3530 0.8808 0.3531 0.5006 1.0454 0.7778
192 168 0.4313 0.4275 0.4347 0.4317 0.4281 0.4238 0.4217 0.4285 0.4257 0.4297 0.5061 0.4929 1.5407 1.0065 1.2732 1.0603 1.1051 0.7924
720 168 0.4424 0.4452 0.4534 0.4598 0.4340 0.4425 0.4675 0.4656 0.4259 0.4389 0.6142 0.5751 2.0399 1.1829 0.9371 0.8819 1.2069 0.8604
1440 168 0.4392 0.4478 0.5222 0.5015 0.4828 0.4759 0.4410 0.4607 0.4646 0.4670 0.7646 0.6819 1.6900 1.0899 1.2457 1.0046 1.0785 0.7730
Average 0.4364 0.4363 0.4610 0.4560 0.4418 0.4401 0.4401 0.4463 0.4372 0.4420 0.6046 0.5634 1.6559 1.0400 0.9523 0.8618 1.1090 0.8009
ETTh2 96 168 0.3541 0.3837 0.3687 0.3925 0.3513 0.3816 0.3727 0.3904 0.3782 0.4076 0.4098 0.4232 3.3130 1.4568 2.5597 1.4723 2.1382 1.0701
192 168 0.3476 0.3869 0.3729 0.3987 0.3532 0.3859 0.3639 0.3923 0.3495 0.3863 0.4312 0.4568 3.9213 1.6013 6.0039 2.2822 1.5139 0.9363
720 168 0.3324 0.3811 0.3716 0.4128 0.3661 0.4012 0.3760 0.4012 0.3984 0.4288 0.4313 0.4673 2.0855 1.2501 6.7391 2.1875 1.4087 0.8958
1440 168 0.3231 0.3787 0.3792 0.4178 0.3911 0.4238 0.4093 0.4307 0.4645 0.4705 0.5024 0.5160 3.5703 1.5863 5.9632 1.9766 3.2192 1.3972
Average 0.3393 0.3826 0.3731 0.4055 0.3654 0.3981 0.3805 0.4036 0.3976 0.4233 0.4437 0.4658 3.2225 1.4736 5.3165 1.9796 2.0700 1.0748
ETTm1 96 192 0.3757 0.3846 0.3717 0.3913 0.3699 0.3857 0.3800 0.3926 0.3844 0.3920 0.6755 0.5449 1.6873 1.0646 0.5354 0.6012 0.9690 0.7273
192 192 0.3446 0.3715 0.3566 0.3857 0.3437 0.3753 0.3673 0.3889 0.3513 0.3822 0.4905 0.4949 1.2954 0.8786 0.5323 0.6288 1.1258 0.8013
720 192 0.3485 0.3774 0.3616 0.3908 0.3528 0.3875 0.3379 0.3787 0.3563 0.3888 0.6736 0.5495 2.1222 1.2104 0.8171 0.8229 1.0007 0.7421
1440 192 0.3438 0.3789 0.4025 0.4256 0.4029 0.4210 0.3653 0.3960 0.3904 0.4242 0.6800 0.5600 1.6025 0.9769 0.6516 0.6802 1.2127 0.8443
average 0.3532 0.3781 0.3731 0.3983 0.3673 0.3924 0.3626 0.3890 0.3706 0.3968 0.6299 0.5373 1.6769 1.0326 0.6341 0.6833 1.0770 0.7788
ETTm2 96 192 0.2445 0.3043 0.2486 0.3099 0.2420 0.3019 0.2582 0.3201 0.2453 0.3133 0.2863 0.3406 3.4978 1.5438 0.5354 0.6012 1.4434 0.8721
192 192 0.2296 0.2959 0.2579 0.3146 0.2393 0.2988 0.2461 0.3089 0.2451 0.3094 0.2986 0.3539 2.3569 1.2538 1.4877 0.9955 1.5398 0.9198
720 192 0.2200 0.2984 0.2962 0.3344 0.2702 0.3262 0.2353 0.3071 0.2350 0.3137 0.3426 0.3856 2.7441 1.2724 2.6198 1.4873 1.6616 0.9615
1440 192 0.2269 0.3048 0.2545 0.3315 0.2467 0.3163 0.2492 0.3157 0.2738 0.3531 0.3761 0.4220 3.1531 1.4227 11.7576 3.2698 1.7756 1.0054
average 0.2303 0.3009 0.2643 0.3226 0.2495 0.3108 0.2472 0.3129 0.2498 0.3224 0.3259 0.3259 2.9380 1.3732 4.1001 1.5884 1.6051 0.9397
electricity 96 168 0.1883 0.2730 0.1665 0.2567 0.1748 0.2625 0.1679 0.2592 0.1893 0.2763 0.5852 0.5886 1.4058 0.9506 0.6377 0.6256 —— ——
192 168 0.1612 0.2533 0.1503 0.2425 0.1536 0.2452 0.1511 0.2440 0.1622 0.2567 0.2747 0.3772 1.5622 1.0112 1.0692 0.8220 —— ——
720 168 0.1466 0.2414 0.1517 0.2481 0.1481 0.2464 0.1491 0.2479 0.1501 0.2493 0.3037 0.3932 1.4369 1.0072 4.0640 1.4426 —— ——
1440 168 0.1463 0.2428 0.1553 0.2546 0.1486 0.2446 0.1466 0.2465 0.1495 0.2501 0.4864 0.5092 1.5501 1.0050 1.0536 0.8189 —— ——
average 0.1606 0.2526 0.1559 0.2505 0.1563 0.2497 0.1537 0.2494 0.1628 0.2581 0.4125 0.4671 1.4888 0.9935 1.7061 0.9273 —— ——
Exchange 96 14 0.0162 0.0830 0.0169 0.0834 0.0158 0.0807 0.0156 0.0825 0.0156 0.0799 0.0847 0.2084 2.5775 1.3205 3.1994 1.4322 1.1299 0.7540
192 14 0.0180 0.0892 0.0182 0.0885 0.0174 0.0849 0.0171 0.0883 0.0166 0.0838 0.1450 0.2834 3.1525 1.4267 7.1367 2.9513 1.1609 0.7874
720 14 0.0204 0.0992 0.0206 0.0999 0.0205 0.0964 0.0521 0.1683 0.0236 0.1092 0.6926 0.6549 2.2854 1.2949 1.0775 0.8829 1.4166 0.9583
1440 14 0.0269 0.1202 0.1166 0.1882 0.0300 0.1203 0.0324 0.1320 0.1091 0.2550 1.4616 0.9538 2.8615 1.3872 7.9106 2.7104 1.5019 0.9877
average 0.0204 0.0979 0.0430 0.1150 0.0209 0.0955 0.0293 0.1178 0.0412 0.1320 0.5960 0.5960 2.7192 1.3573 4.8311 1.9942 1.3023 0.8719
Weather 96 672 0.3529 0.3470 0.3432 0.3420 0.3383 0.3384 0.3553 0.3490 0.3585 0.3988 0.3770 0.3742 —— —— —— —— 0.4796 0.4392
192 672 0.3401 0.3403 0.3285 0.3451 0.3245 0.3335 0.3398 0.3409 0.3502 0.3962 0.3720 0.3800 —— —— —— —— 0.5108 0.4526
720 672 0.3142 0.3314 0.3337 0.3426 0.3200 0.3404 0.3179 0.3383 0.3266 0.3659 0.4117 0.4310 —— —— —— —— 0.4697 0.4512
1440 672 0.3097 0.3304 0.3126 0.3349 0.3437 0.3596 0.3067 0.3318 0.3237 0.3595 0.3798 0.4050 —— —— —— —— 0.4765 0.4561
average 0.3292 0.3373 0.3295 0.3411 0.3316 0.3430 0.3299 0.3400 0.3398 0.3801 0.3851 0.3976 —— —— —— —— 0.4842 0.4498
Wind 96 192 1.1953 0.7605 1.1628 0.7446 1.1471 0.7426 1.1796 0.7621 1.0439 0.7215 1.3067 0.8154 2.1130 1.0870 3.7312 1.3789 1.3909 0.8543
192 192 1.1693 0.7601 1.1904 0.7582 1.1597 0.7477 1.1637 0.7639 1.0481 0.7211 1.3662 0.8590 1.8822 1.0471 3.7236 1.4758 1.4510 0.8885
720 192 1.0969 0.7605 1.1576 0.7643 1.2393 0.7897 1.0970 0.7569 1.1007 0.7498 1.3665 0.8780 1.9993 1.1442 3.8525 1.4478 1.5825 0.9126
1440 192 1.0434 0.7326 1.0880 0.7658 1.4168 0.8580 1.1244 0.7597 1.1375 0.7587 1.2493 0.8433 2.2880 1.1957 3.4273 1.3756 1.5407 0.9212
average 1.1262 0.7534 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 1.3222 0.8489 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
The experimental results in Table 6 provide strong evidence of the Flexible Decoupled Framework’s (FDF) effectiveness in time series forecasting, consistently outperforming state-of-the-art models in both Mean Squared Error (MSE) and Mean Absolute Error (MAE) across diverse datasets. This advantage stems from FDF’s decoupling strategy, which separates trend and seasonal components for more precise modeling of temporal patterns. Notably, FDF excels on challenging datasets like ETTh1 and ETTh2, where traditional models like Diffusion-TS and D3VAE struggle with varying time resolutions and high noise levels. FDF’s resilience, attributed to the Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), enables it to capture complex trends and seasonal dynamics with top performance in both MSE and MAE. In high-dimensional datasets like Electricity and Weather, FDF maintains superior accuracy, showcasing its robustness in handling intricate temporal dependencies. Unlike models such as PatchMixer and Diffusion-TS, which face difficulties with long forecast horizons, FDF effectively manages long-term dependencies. On more volatile datasets like Wind, FDF captures sharp fluctuations and demonstrates competitive results, though further refinement could enhance performance on extremely noisy data. Overall, FDF demonstrates strong generalization across datasets, managing both short- and long-term patterns better than methods like DLinear and Autoformer. In summary, FDF’s decoupling significantly enhances forecasting performance, allowing it to outperform leading diffusion-based and traditional models. This validates the framework’s theoretical contributions, where decoupling was hypothesized to improve accuracy and interpretability, now confirmed by the experimental results.
C.2. Decoupling Variants Experiments
As shown in table 7, the evaluation compares two model variants:
  • FDFw/o ${}_{\text{w/o decoupling}}$: This variant bypasses the decoupling process, directly applying the CDSM to the original time series data.
[…]
Table 7. Performance comparison of decoupling methods on ETTh2 and ETTm2.
[…]
MSE MAE MSE MAE MSE MAE
ETTh2 0.3541 0.3837 0.3609 0.3874 0.3712 0.3947
0.3476 0.3869 0.4039 0.4174 0.3878 0.4069
0.3324 0.3811 0.4015 0.4258 0.3904 0.4153
0.3231 0.3787 0.3691 0.4118 0.3801 0.4139
Average 0.3393 0.3826 0.3839 0.4106 0.3824 0.4077
ETTm2 0.2446 0.3034 0.2435 0.3039 0.2456 0.3057
0.2315 0.2956 0.2396 0.3013 0.2502 0.3102
0.2175 0.2920 0.2488 0.3159 0.2517 0.3246
0.2349 0.3175 0.2632 0.3383 0.2737 0.3418
Average 0.2321 0.3021 0.2488 0.3148 0.2553 0.3206
Table 8. Performance comparison across different variants on ETTh2 and ETTm2.
Dataset L T FDF FDFCR FDFPLR FDF3 FDF4 FDF5
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 96 192 0.3541 0.3837 0.3690 0.3889 0.3518 0.3819 0.3794 0.4138 0.3764 0.3940 0.3650 0.3873
192 192 0.3476 0.3869 0.3480 0.3801 0.3466 0.3822 0.3731 0.4084 0.3426 0.3745 0.3403 0.3811
720 192 0.3324 0.3811 0.3287 0.3732 0.3395 0.3806 0.3800 0.4212 0.3444 0.3782 0.3314 0.3756
1440 192 0.3231 0.3787 0.3158 0.3701 0.3354 0.3856 0.4088 0.4427 0.3529 0.3862 0.3254 0.3819
Average 0.3393 0.3826 0.3404 0.3781 0.3433 0.3826 0.3853 0.4215 0.3541 0.3832 0.3405 0.3815
ETTm2 96 192 0.2445 0.3043 0.2461 0.3042 0.2453 0.3049 0.2584 0.3370 0.2585 0.3370 0.2465 0.3043
192 192 0.2296 0.2959 0.2307 0.2953 0.2307 0.2970 0.2636 0.3411 0.2642 0.3417 0.2297 0.2961
720 192 0.2200 0.2984 0.2279 0.3036 0.2379 0.3108 0.2323 0.3138 0.2331 0.3146 0.2192 0.2986
1440 192 0.2269 0.3048 0.2232 0.3061 0.2659 0.3356 0.2293 0.3170 0.2295 0.3172 0.2268 0.3049
Average 0.2303 0.3009 0.2320 0.3023 0.2450 0.3121 0.2459 0.3272 0.2463 0.3276 0.2305 0.3010
C.3. Ablation experiments
We conduct a comprehensive ablation study to evaluate the impact of key components in the proposed FDF architecture. Specifically, we examine the role of the condition information injection module, the PTM, and instance normalization in handling the complex temporal dynamics present in time series data. The study is performed on the ETTh2 and ETTm2 datasets, and the results are presented in Table 8.
  • FDFw/o ${}_{\text{w/o CR}}$: Removes the conditional information.
[…]
  • FDFw/o ${}_{\text{w/o INR}}$: Removes instance normalization during both the training and inference phases.
  • FDFw/o ${}_{\text{w/o CR+PLR}}$: Removes the conditional information and replaces PTM with a linear layer.
  • FDFw/o ${}_{\text{w/o INA}}$: Removes instance normalization alignment.
The results show that each component in FDF impacts overall performance to varying degrees, particularly in modeling complex temporal patterns. Removing the condition information in FDFCR leads to a slight increase in MSE on ETTh2 (from 0.3393 to 0.3404), indicating that while this module helps integrate external information, the model remains resilient without it, though with reduced flexibility. More significant degradation occurs when replacing the PTM with a simple linear layer (FDFPLR), especially on ETTm2, where MSE increases from 0.2303 to 0.2450, emphasizing the importance of the non-linear trend modeling provided by the PTM. The removal of instance normalization (FDF3) results in the sharpest performance drop, particularly on ETTm2, where MSE rises to 0.2459. This underscores the critical role of instance normalization in stabilizing training and preventing overfitting on datasets with significant temporal variations. Removing both the condition information and the PTM in FDF4 leads to further performance degradation, showing that these components play complementary roles in improving model performance. Finally, eliminating instance normalization alignment during training (FDF5) yields mixed results. On ETTh2, a slight improvement suggests that aligning normalization parameters may hinder the model’s ability to adapt to complex dynamics, while on ETTm2, where temporal dependencies dominate, the effect is minimal. Overall, the ablation study validates FDF’s modular design in improving adaptability and accuracy across diverse time series forecasting tasks.
Figure 8. Comparison of FDF and deterministic models on ETTh1 Dataset.
Figure 9. Comparison of FDF and deterministic models on ETTh2 Dataset.
C.4. Joint Modeling Capability Visualization
The visualizations in Figures 10 and 11 illustrate the interpretability and effectiveness of the proposed FDF framework. By comparing the predicted and actual trend and seasonal components, these figures highlight FDF’s capacity to handle diverse temporal structures, validating its robustness across different time series conditions. In Figure 10, which depicts the Electricity dataset known for its clear periodicity and pronounced trends, FDF demonstrates strong predictive capabilities. The Polynomial Trend Module (PTM) captures smooth, large-scale patterns, evident from the close alignment between the predicted (red) and true (green) trend lines in subfigure (a). This precise trend prediction highlights PTM’s ability to model broad temporal dynamics effectively. Similarly, the Conditional Denoising Seasonal Module (CDSM) accurately models complex periodic fluctuations, as shown in subfigure (b), where the predicted and actual seasonal components closely match. These results emphasize the effectiveness of FDF’s decoupled approach, particularly its diffusion-based mechanism for handling distinct temporal features with high fidelity. In contrast, Figure 11, representing the more irregular and volatile ETTh1 dataset, demonstrates FDF’s adaptability to complex temporal dynamics. The PTM captures broad underlying patterns despite data volatility, as seen in subfigure (a). In subfigure (b), the CDSM effectively models sharp, frequent seasonal fluctuations, further illustrating FDF’s versatility in handling irregular temporal patterns. Overall, these visualizations confirm FDF’s ability to strike a balance between preserving trends and accurately capturing variations. The model performs well on the Electricity dataset, where strong trend alignment is observed, and shows resilience on the more volatile ETTh1 dataset. These results not only affirm the interpretability and robustness of the model but also highlight its adaptability in addressing diverse forecasting challenges. FDF’s consistent performance across different datasets reinforces its potential as a reliable and generalizable solution for complex time series forecasting tasks.
Figure 10. Visualization of trend and seasonal components for the electricity dataset.
Figure 11. Visualization of trend and seasonal components for the ETTh1 dataset.
C.5. Case Study
In this case study, we systematically compare the performance of FDF against several deterministic and diffusion-based models on the ETTh1 and ETTh2 datasets, as shown in Figures 8 and 9. With a history length of 192, FDF consistently outperforms other models in capturing both trends and seasonal variations. On the ETTh1 dataset, FDF aligns closely with the ground truth, excelling in modeling complex seasonal fluctuations and major trend shifts. In contrast, deterministic models such as PatchMixer and TSMixer struggle to capture rapid trend changes and local seasonal variations, leading to less accurate predictions. Similarly, Autoformer and DLinear, while effective at trend modeling, often produce delayed or overly smooth predictions, missing key variations. FDF, by striking an optimal balance between trend and seasonal dynamics, better handles these complex time series patterns. Figures 12 and Figures13 compare FDF with diffusion-based models like CSDI and Diffusion-TS. FDF outperforms these models, delivering smoother and more accurate forecasts. Diffusion-TS, in particular, exhibits higher variability and less reliable predictions, especially on the ETTh2 dataset, which contains intricate seasonal and trend components. FDF avoids over-smoothing and effectively manages complex temporal structures, demonstrating superior resilience in such scenarios. This study highlights FDF’s effectiveness in capturing both trends and fluctuations in complex datasets like ETTh1 and ETTh2. Compared to deterministic models, which either oversmooth predictions or struggle with rapid transitions, FDF consistently achieves a better balance in modeling diverse temporal dynamics. Additionally, FDF outperforms diffusion-based models by providing more reliable forecasts, even on datasets with intricate multi-scale patterns. These results underscore the advantages of FDF’s decoupling strategy and specialized modules, positioning it as a superior solution for tasks involving complex, multi-scale temporal dynamics.
Figure 12. Comparison of FDF and DDPM Models on ETTh1 Dataset.
Figure 13. Comparison of FDF and DDPM Models on ETTh2 Dataset.